# PersoBench: Benchmarking Personalized Response Generation in Large Language Models

### v1

@@ 6 CONCLUSION AND FUTURE WORK
In this paper, we introduced PersoBench, an automatic benchmarking pipeline designed to evaluate LLMs in the context of persona-aware conversation generation. Using both vanilla and COT prompting setups in a zero-shot setting, we benchmarked well-known open-source and closed-source LLMs on three persona-aware datasets. Leveraging eight well-established evaluation metrics, we assessed the models across fluency, diversity, coherence, and personalization dimensions. Our findings suggest that while current LLMs perform well in generating fluent and diverse responses, they struggle to deliver personalized and coherent responses in persona-augmented conversational contexts.
A potential future direction for research in this domain could involve exploring beyond textual persona representation to include other forms, such as tabular demographic data or multi-modal setups incorporating different media types.
@@
While significant effort has been invested in this research, there are still some limitations to our study: (1) In crafting an automatic evaluation pipeline, we required the LLMs to format their generated output in JSON format to enable the automatic extraction of the response and reasoning. However, this automation discards responses that could not be successfully parsed as JSON, necessitating further manual intervention and potentially extending the post-processing workload with human resources. (2) To increase the likelihood of successful response generation in JSON format, we opted for instruct-tuned versions of LLMs. While this choice was necessary for the study, it excludes the evaluation of standard versions of LLMs that may perform well in more general, unstructured conversational contexts.
Acknowledgments
We acknowledge the Centre for Applied Artificial Intelligence at Macquarie University (Sydney, Australia) for funding this research.
@@ Acknowledgments
Table 9: Explicit Prompt Examples.
@@
I will provide you with a conversation context and the personas of the participants, that can be annotated with speaker information. As a participant in this conversation, your task is to generate a personalized response, considering the conversation context and the annotated personas. Participant Personas: [User 1 persona]: [’i attend book club every week.’ ’i am more of a cat person than a dog person.’] Conversation Context: User1: my parents dies in an accident User2: that is sad . my best friend is my mom. i could not image loosing her. User1: it is harder some days than others. but it is good you get along with your mom so well. Task Instruction: * Provide an unannotated response. * If only one persona is available, personalize the response accordingly. * If the conversation context is a single query, respond appropriately to the query. * Apply Chain of Thought reasoning to reflect on the alignment of your response with the personas. Output Format: Output a JSON of the following format: { "reasoning": "briefly describe your personalization process (in 110 words or less)." "response": "provide the personalized natural language response here." }
Table 10: Sample COT setup prompt.
7.2 Metrics
To evaluate fluency, we used ROUGE [26], METEOR [27], and BERTScore [28]. ROUGE, via the rouge-score package [^8] ROUGE-Score package: https://github.com/google-research/google-research/tree/master/rouge, measures n-gram overlap between generated and reference texts. METEOR, implemented through NLTK [^9] NLTK package: https://www.nltk.org/, accounts for synonymy and stemming, offering a more nuanced linguistic evaluation. BERTScore, implemented using the bert-score package [^10] BERTScore package: https://github.com/Tiiiger/bert_score, evaluates Precision, Recall, and F1 based on contextual embeddings, assessing how well the generated text preserves meaning and coherence relative to the reference.
@@ 7.2 Metrics
For diversity assessment, we calculated Dist-1 and Dist-2 metrics [29], which respectively measure the ratio of unique unigrams and bigrams in the generated outputs.
Additionally, We used advanced metrics to evaluate personalization and coherence in conversational agents. Personalization can be evaluated from two aspects: consistency and coverage. Consistency reflects whether the generated response R is aligned with the given personal information P, while coverage indicates how much of the given personal information is reflected in the generated response. The C Score [30], a metric for persona consistency, evaluates whether the responses remain consistent with the persona by fine-tuning BERT on Dialogue Natural Language Inference (DNLI).
On the other hand, the Persona Distance [31], a metric for persona coverage, measures the semantic alignment between the response and persona by utilizing word2vec cosine similarity. Additionally, the Utterance Entailment Score (UE-Score) [32] measures dialogue coherence by calculating the NLI alignment score between an utterance and its response. The UE-Score model is based on BERT, which has been fine-tuned using the Stanford Natural Language Inference (SNLI) [^11] https://nlp.stanford.edu/projects/snli/ dataset [34]. These metrics are essential for a nuanced assessment of personalized dialogue systems.
Table 11: Fluency evaluation of LLMs on the IT-ConvAI2 dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.12 $\pm$ 0.1 / 0.02 $\pm$ 0.05 / 0.11 $\pm$ 0.09 0.14 $\pm$ 0.11 0.81 $\pm$ 0.18 / 0.81 $\pm$ 0.18 / 0.81 $\pm$ 0.18 0.09 $\pm$ 0.08 / 0.01 $\pm$ 0.04 / 0.09 $\pm$ 0.07 0.11 $\pm$ 0.09 0.80 $\pm$ 0.18 / 0.80 $\pm$ 0.18 / 0.80 $\pm$ 0.18
Qwen2 7B 0.10 $\pm$ 0.08 / 0.02 $\pm$ 0.04 / 0.10 $\pm$ 0.08 0.14 $\pm$ 0.1 0.83 $\pm$ 0.12 / 0.84 $\pm$ 0.12 / 0.83 $\pm$ 0.12 0.09 $\pm$ 0.06 / 0.01 $\pm$ 0.03 / 0.08 $\pm$ 0.06 0.12 $\pm$ 0.09 0.82 $\pm$ 0.08 / 0.83 $\pm$ 0.08 / 0.83 $\pm$ 0.08
Llama3.1 8B 0.11 $\pm$ 0.09 / 0.02 $\pm$ 0.05 / 0.10 $\pm$ 0.09 0.13 $\pm$ 0.10 0.76 $\pm$ 0.25 / 0.77 $\pm$ 0.25 / 0.76 $\pm$ 0.25 0.11 $\pm$ 0.09 / 0.02 $\pm$ 0.04 / 0.10 $\pm$ 0.08 0.12 $\pm$ 0.09 0.84 $\pm$ 0.07 / 0.83 $\pm$ 0.07 / 0.84 $\pm$ 0.07
GPT3.5 Turbo 0.12 $\pm$ 0.09 / 0.02 $\pm$ 0.05 / 0.11 $\pm$ 0.08 0.15 $\pm$ 0.11 0.85 $\pm$ 0.02 / 0.85 $\pm$ 0.03/ 0.85 $\pm$ 0.02 0.11 $\pm$ 0.08 / 0.02 $\pm$ 0.04 / 0.10 $\pm$ 0.07 0.13 $\pm$ 0.10 0.84 $\pm$ 0.03 / 0.84 $\pm$ 0.03 / 0.84 $\pm$ 0.03
GPT4o Mini 0.11 $\pm$ 0.07 / 0.02 $\pm$ 0.04 / 0.10 $\pm$ 0.07 0.16 $\pm$ 0.10 0.84 $\pm$ 0.05 / 0.85 $\pm$ 0.05 / 0.84 $\pm$ 0.05 0.10 $\pm$ 0.06 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.06 0.16 $\pm$ 0.09 0.83 $\pm$ 0.04 / 0.85 $\pm$ 0.04 / 0.84 $\pm$ 0.04
GPT4 Turbo 0.12 $\pm$ 0.09 / 0.02 $\pm$ 0.05 / 0.11 $\pm$ 0.08 0.15 $\pm$ 0.11 0.85 $\pm$ 0.02 / 0.86 $\pm$ 0.03 / 0.85 $\pm$ 0.02 0.11 $\pm$ 0.08 / 0.02 $\pm$ 0.04 / 0.10 $\pm$ 0.08 0.13 $\pm$ 0.10 0.84 $\pm$ 0.02 / 0.84 $\pm$ 0.02 / 0.84 $\pm$ 0.02
Table 12: Fluency evaluation of LLMs on the Blended Skill Talk dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.10 $\pm$ 0.08 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.07 0.11 $\pm$ 0.08 0.80 $\pm$ 0.19 / 0.81 $\pm$ 0.19 / 0.81 $\pm$ 0.19 0.10 $\pm$ 0.07 / 0.01 $\pm$ 0.02 / 0.09 $\pm$ 0.07 0.11 $\pm$ 0.08 0.80 $\pm$ 0.18 / 0.82 $\pm$ 0.18 / 0.81 $\pm$ 0.18
Qwen2 7B 0.11 $\pm$ 0.07 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.07 0.12 $\pm$ 0.08 0.82 $\pm$ 0.13 / 0.84 $\pm$ 0.14 / 0.83 $\pm$ 0.13 0.09 $\pm$ 0.06 / 0.01 $\pm$ 0.02 / 0.09 $\pm$ 0.06 0.11 $\pm$ 0.07 0.83 $\pm$ 0.05 / 0.86 $\pm$ 0.05 / 0.84 $\pm$ 0.05
Llama3.1 8B 0.12 $\pm$ 0.08 / 0.02 $\pm$ 0.04 / 0.11 $\pm$ 0.08 0.12 $\pm$ 0.09 0.81 $\pm$ 0.17 / 0.82 $\pm$ 0.18 / 0.82 $\pm$ 0.17 0.11 $\pm$ 0.08 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.07 0.11 $\pm$ 0.08 0.85 $\pm$ 0.02 / 0.86 $\pm$ 0.02 / 0.85 $\pm$ 0.02
GPT3.5 Turbo 0.13 $\pm$ 0.09 / 0.02 $\pm$ 0.04 / 0.11 $\pm$ 0.08 0.13 $\pm$ 0.09 0.85 $\pm$ 0.02 / 0.86 $\pm$ 0.02 / 0.86 $\pm$ 0.02 0.11 $\pm$ 0.07 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.06 0.12 $\pm$ 0.08 0.82 $\pm$ 0.13 / 0.84 $\pm$ 0.13 / 0.83 $\pm$ 0.13
GPT4o Mini 0.13 $\pm$ 0.08 / 0.02 $\pm$ 0.03 / 0.12 $\pm$ 0.07 0.14 $\pm$ 0.09 0.85 $\pm$ 0.03 / 0.87 $\pm$ 0.04 / 0.86 $\pm$ 0.03 0.12 $\pm$ 0.07 / 0.01 $\pm$ 0.03 / 0.11 $\pm$ 0.06 0.14 $\pm$ 0.08 0.84 $\pm$ 0.02 / 0.86 $\pm$ 0.02 / 0.85 $\pm$ 0.02
GPT4 Turbo 0.12 $\pm$ 0.08 / 0.02 $\pm$ 0.04 / 0.11 $\pm$ 0.07 0.12 $\pm$ 0.09 0.85 $\pm$ 0.02 / 0.86 $\pm$ 0.02 / 0.86 $\pm$ 0.02 0.11 $\pm$ 0.07 / 0.01 $\pm$ 0.03 / 0.10 $\pm$ 0.07 0.12 $\pm$ 0.08 0.84 $\pm$ 0.03 / 0.86 $\pm$ 0.04 / 0.85 $\pm$ 0.03
Table 13: Fluency evaluation of LLMs on the FoCus dataset.
Model Vanilla Setup COT Setup
ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1) ROUGE (R1/R2/RL) METEOR BERTScore (Prec/Rec/F1)
Mistral 7B 0.11 $\pm$ 0.12 / 0.03 $\pm$ 0.06 / 0.10 $\pm$ 0.11 0.11 $\pm$ 0.14 0.48 $\pm$ 0.42 / 0.49 $\pm$ 0.43 / 0.48 $\pm$ 0.42 0.15 $\pm$ 0.09 / 0.03 $\pm$ 0.04 / 0.13 $\pm$ 0.08 0.15 $\pm$ 0.10 0.75 $\pm$ 0.25 / 0.77 $\pm$ 0.26 / 0.76 $\pm$ 0.26
Qwen2 7B 0.08 $\pm$ 0.10 / 0.02 $\pm$ 0.04 / 0.08 $\pm$ 0.09 0.10 $\pm$ 0.13 0.41 $\pm$ 0.42 / 0.43 $\pm$ 0.44 / 0.42 $\pm$ 0.43 0.10 $\pm$ 0.08 / 0.02 $\pm$ 0.03 / 0.09 $\pm$ 0.07 0.13 $\pm$ 0.10 0.60 $\pm$ 0.37 / 0.63 $\pm$ 0.38 / 0.62 $\pm$ 0.38
Llama3.1 8B 0.11 $\pm$ 0.15 / 0.03 $\pm$ 0.1 / 0.10 $\pm$ 0.14 0.11 $\pm$ 0.14 0.47 $\pm$ 0.42 / 0.47 $\pm$ 0.43 / 0.47 $\pm$ 0.43 0.15 $\pm$ 0.10 / 0.03 $\pm$ 0.07 / 0.13 $\pm$ 0.09 0.16 $\pm$ 0.10 0.80 $\pm$ 0.18 / 0.81 $\pm$ 0.19 / 0.80 $\pm$ 0.19
GPT3.5 Turbo 0.20 $\pm$ 0.14 / 0.06 $\pm$ 0.12 / 0.18 $\pm$ 0.14 0.20 $\pm$ 0.15 0.78 $\pm$ 0.25 / 0.79 $\pm$ 0.25 / 0.78 $\pm$ 0.25 0.17 $\pm$ 0.10 / 0.04 $\pm$ 0.08 / 0.15 $\pm$ 0.10 0.18 $\pm$ 0.12 0.79 $\pm$ 0.21 / 0.80 $\pm$ 0.22 / 0.80 $\pm$ 0.21
GPT4o Mini 0.19 $\pm$ 0.11 / 0.06 $\pm$ 0.08 / 0.17 $\pm$ 0.10 0.22 $\pm$ 0.15 0.78 $\pm$ 0.24 / 0.81 $\pm$ 0.25 / 0.80 $\pm$ 0.24 0.20 $\pm$ 0.09 / 0.05 $\pm$ 0.07 / 0.17 $\pm$ 0.08 0.22 $\pm$ 0.12 0.85 $\pm$ 0.05 / 0.87 $\pm$ 0.05 / 0.86 $\pm$ 0.05
GPT4 Turbo 0.23 $\pm$ 0.13 / 0.08 $\pm$ 0.11 / 0.20 $\pm$ 0.13 0.23 $\pm$ 0.15 0.84 $\pm$ 0.16 / 0.84 $\pm$ 0.16 / 0.84 $\pm$ 0.16 0.22 $\pm$ 0.11 / 0.07 $\pm$ 0.09 / 0.19 $\pm$ 0.10 0.23 $\pm$ 0.14 0.85 $\pm$ 0.10 / 0.86 $\pm$ 0.10 / 0.86 $\pm$ 0.11
Table 14: Diversity evaluation of LLMs using Dist-1 and Dist-2 metrics across benchmarking datasets in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
Mistral 7B 0.86 $\pm$ 0.2 0.95 $\pm$ 0.21 0.87 $\pm$ 0.2 0.95 $\pm$ 0.21 0.87 $\pm$ 0.15 0.94 $\pm$ 0.22 0.86 $\pm$ 0.2 0.95 $\pm$ 0.21 0.49 $\pm$ 0.44 0.56 $\pm$ 0.49 0.75 $\pm$ 0.26 0.88 $\pm$ 0.3
Qwen2 7B 0.90 $\pm$ 0.14 0.98 $\pm$ 0.14 0.90 $\pm$ 0.1 0.99 $\pm$ 0.09 0.90 $\pm$ 0.15 0.97 $\pm$ 0.15 0.91 $\pm$ 0.07 0.99 $\pm$ 0.06 0.43 $\pm$ 0.44 0.49 $\pm$ 0.5 0.60 $\pm$ 0.37 0.72 $\pm$ 0.44
Llama3.1 8B 0.78 $\pm$ 0.27 0.89 $\pm$ 0.3 0.89 $\pm$ 0.1 0.99 $\pm$ 0.08 0.84 $\pm$ 0.19 0.95 $\pm$ 0.2 0.90 $\pm$ 0.07 0.99 $\pm$ 0.04 0.44 $\pm$ 0.41 0.53 $\pm$ 0.48 0.76 $\pm$ 0.19 0.93 $\pm$ 0.21
GPT3.5 Turbo 0.91 $\pm$ 0.06 1.0 $\pm$ 0.01 0.91 $\pm$ 0.06 0.99 $\pm$ 0.03 0.92 $\pm$ 0.06 1.0 $\pm$ 0.01 0.87 $\pm$ 0.15 0.97 $\pm$ 0.15 0.79 $\pm$ 0.26 0.90 $\pm$ 0.28 0.78 $\pm$ 0.22 0.92 $\pm$ 0.25
GPT4o Mini 0.89 $\pm$ 0.07 0.99 $\pm$ 0.05 0.88 $\pm$ 0.06 0.99 $\pm$ 0.04 0.91 $\pm$ 0.06 1.0 $\pm$ 0.03 0.89 $\pm$ 0.05 1.0 $\pm$ 0.01 0.79 $\pm$ 0.25 0.91 $\pm$ 0.28 0.85 $\pm$ 0.07 0.99 $\pm$ 0.06
GPT4 Turbo 0.93 $\pm$ 0.05 1.0 $\pm$ 0.01 0.93 $\pm$ 0.05 1.0 $\pm$ 0.01 0.92 $\pm$ 0.05 1.0 $\pm$ 0.01 0.91 $\pm$ 0.06 1.0 $\pm$ 0.03 0.86 $\pm$ 0.18 0.96 $\pm$ 0.18 0.86 $\pm$ 0.11 0.98 $\pm$ 0.11
@@
Each of these metrics offers distinct insights into the model’s performance, collectively ensuring a comprehensive and nuanced evaluation of the generated responses. When the model’s output couldn’t be evaluated because of parsing errors, we assigned default worst-case values. While discarding unparsable outputs, it is essential to ensure the automation of our benchmark.
7.3 Comprenehsive Results
Tables 11-17 present the detailed experimental results for all evaluated aspects, including both the average metric values and their standard deviations.
Table 15: Coherence evaluation using UE Score on benchmarking datasets in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
Mistral 7B 0.46 ± 0.72 0.31 ± 0.67 0.48 ± 0.79 0.44 ± 0.79 0.21 ± 0.52 0.33 ± 0.62
Qwen2 7B 0.42 ± 0.73 0.38 ± 0.72 0.39 ± 0.75 0.52 ± 0.84 0.17 ± 0.48 0.36 ± 0.70
Llama3.1 8B 0.23 ± 0.58 0.22 ± 0.58 0.33 ± 0.70 0.37 ± 0.73 0.17 ± 0.48 0.33 ± 0.64
GPT 3.5 Turbo 0.49 ± 0.76 0.38 ± 0.74 0.51 ± 0.82 0.57 ± 0.86 0.37 ± 0.67 0.45 ± 0.72
GPT 4o Mini 0.40 ± 0.75 0.38 ± 0.74 0.38 ± 0.74 0.39 ± 0.76 0.35 ± 0.69 0.37 ± 0.70
GPT4 Turbo 0.43 ± 0.78 0.42 ± 0.78 0.56 ± 0.84 0.50 ± 0.82 0.39 ± 0.73 0.45 ± 0.76
Table 16: Personalization evaluation of LLMs using P-Dist Score and C Score across the benchmarking dataset, in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
P-Dist Score ±SD C Score ±SD P-Dist Score ±SD C Score ±SD P-Dist Score ±SD C Score ±SD P-Dist Score ±SD C Score ±SD P-Dist Score ±SD C Score ±SD P-Dist Score ±SD C Score ±SD
Mistral-7B 0.62 ±0.20 0.33 ±0.73 0.54 ±0.20 0.02 ±0.67 0.46 ±0.18 -0.00 ±0.59 0.43 ±0.17 -0.08 ±0.59 0.30 ±0.28 -0.23 ±0.82 0.53 ±0.22 0.41 ±0.75
Qwen2-7B 0.61 ±0.17 0.21 ±0.70 0.63 ±0.13 0.23 ±0.67 0.46 ±0.15 -0.11 ±0.61 0.49 ±0.13 0.01 ±0.57 0.29 ±0.30 -0.31 ±0.84 0.43 ±0.28 0.21 ±0.89
Llama3.1-8B 0.55 ±0.22 0.03 ±0.72 0.58 ±0.13 0.05 ±0.63 0.45 ±0.17 -0.11 ±0.61 0.46 ±0.14 -0.05 ±0.53 0.33 ±0.33 -0.22 ±0.87 0.59 ±0.19 0.54 ±0.68
GPT3.5-Turbo 0.63 ±0.15 0.23 ±0.69 0.60 ±0.13 0.10 ±0.62 0.47 ±0.15 -0.01 ±0.60 0.48 ±0.16 0.02 ±0.60 0.51 ±0.21 0.26 ±0.75 0.52 ±0.18 0.40 ±0.72
GPT-40-Mini 0.63 ±0.12 0.10 ±0.63 0.64 ±0.11 0.14 ±0.61 0.48 ±0.13 -0.10 ±0.52 0.50 ±0.13 -0.05 ±0.52 0.54 ±0.19 0.16 ±0.71 0.57 ±0.11 0.30 ±0.63
GPT-4-Turbo 0.60 ±0.13 0.08 ±0.62 0.59 ±0.12 0.01 ±0.58 0.47 ±0.14 -0.05 ±0.53 0.47 ±0.13 -0.05 ±0.53 0.51 ±0.15 0.14 ±0.66 0.52 ±0.13 0.21 ±0.63
Table 17: Response Time and Failure Ratio of LLMs on IT-ConvAI2, Blended Skill Talk, and FoCus datasets, in Vanilla and COT setups.
Model IT-ConvAI2 Blended Skill Talk FoCus
Vanilla COT Vanilla COT Vanilla COT
Resp. Time Fail Ratio Resp. Time Fail Ratio Resp. Time Fail Ratio Resp. Time Fail Ratio Resp. Time Fail Ratio Resp. Time Fail Ratio
Mistral 7B 3.11 ± 1.11 0.05 ± 0.00 6.90 ± 1.61 0.05 ± 0.00 3.87 ± 0.72 0.05 ± 0.00 6.71 ± 1.85 0.05 ± 0.00 4.39 ± 0.36 0.44 ± 0.00 7.67 ± 1.27 0.10 ± 0.00
Qwen2 7B 1.97 ± 0.72 0.02 ± 0.00 4.70 ± 0.94 0.01 ± 0.00 2.21 ± 0.67 0.02 ± 0.00 4.8 ± 0.88 0.00 ± 0.00 3.74 ± 0.76 0.51 ± 0.00 7.28 ± 1.07 0.27 ± 0.00
Llama3.1 8B 4.15 ± 0.53 0.10 ± 0.00 8.17 ± 0.11 0.01 ± 0.00 4.28 ± 0.26 0.04 ± 0.00 8.69 ± 0.15 0.00 ± 0.00 4.54 ± 0.21 0.45 ± 0.00 8.93 ± 0.15 0.05 ± 0.00
GPT3.5 Turbo 1.18 ± 0.57 0.00 ± 0.00 1.40 ± 0.30 0.00 ± 0.00 0.91 ± 0.21 0.00 ± 0.00 1.56 ± 0.47 0.02 ± 0.00 1.30 ± 0.42 0.09 ± 0.00 2.16 ± 0.54 0.07 ± 0.00
GPT4o Mini 0.99 ± 0.37 0.00 ± 0.00 2.20 ± 0.67 0.00 ± 0.00 0.94 ± 0.28 0.00 ± 0.00 1.63 ± 0.42 0.00 ± 0.00 1.43 ± 0.66 0.09 ± 0.00 3.05 ± 8.65 0.00 ± 0.00
GPT4 Turbo 2.21 ± 0.71 0.00 ± 0.00 4.81 ± 1.45 0.00 ± 0.00 1.84 ± 0.5 0.00 ± 0.00 5.15 ± 1.13 0.00 ± 0.00 3.32 ± 1.24 0.04 ± 0.00 7.29 ± 2.43 0.01 ± 0.00

### v2

@@ 7 CONCLUSION AND FUTURE WORK
In this paper, we introduced PersoBench, a new benchmarking pipeline designed to evaluate LLMs in the context of persona-aware conversation generation. Using both vanilla and CoT prompting setups in a zero-shot setting, we benchmarked well-known open-source and closed-source LLMs on three persona-aware datasets. Leveraging eight well-established evaluation metrics, we assessed the models across fluency, diversity, coherence, and personalization dimensions. Our findings suggest that while current LLMs perform well in generating fluent and diverse responses, they struggle to deliver personalized and coherent responses in persona-augmented conversational contexts. A potential future direction for research in this domain could involve exploring beyond textual persona representation to include other forms, such as tabular demographic data or multi-modal setups incorporating different media types.
@@
While significant effort has been invested in this research, several limitations remain. (1) Our automatic evaluation pipeline requires models to produce responses in a strict JSON format to enable reliable parsing of generated responses and reasoning. This constraint led to the exclusion of outputs that failed to conform to the format, occasionally requiring manual post-processing and motivating the use of instruct-tuned only LLM variants. As a result, standard LLMs operating in less structured conversational settings were not evaluated. (2) Due to the high cost of commercial LLM APIs, our experiments were conducted on a limited dataset size, which constrained broader scaling analyses, including more exhaustive comparisons across model configurations and token-length behaviors. A more fine-grained analysis of model performance under varying context lengths is left for future work. (3) Our evaluation primarily relied on automatic metrics. Although we provide qualitative examples for additional insight, a comprehensive human evaluation and metric–human correlation study was beyond the scope of this work and is left to future research.
Appendix
.1 Explicit prompt samples
Gathered from related studies in relevant contexts, Table 8 illustrates explicit samples aligned with the implicit design principles mentioned. It includes prompts from general conversational AI as well as specialized domains such as emotional support, healthcare, and factuality evaluation. Table.9 illustrates one sample of our generated prompts that includes the CoT reasoning.
Table 8: Explicit Prompt Examples.
Prompt Source Prompt Template
Pal Persona-Augmented prompt [12] Persona: {persona}
Dialogue history: {context}
Response:
RoleLLM [40] System Instruction: You are {role_name}, your description is  {role_description_and_catchphrases}. Now please answer some questions to accurately show your personality traits! Your speaking style should fully imitate the personality role assigned to you! Please do not expose that you are an artificial intelligence model or a language model, you must always remember that you are only assigned one personality role. Don’t be verbose or too formal or polite when speaking.
User Prompt:
{user_name}: “{user_instruction}”
FELM Claim Extraction[44] I will show you a question and a list of text segments. The text segments can be concatenated to form a complete answer to the question. Your task is to extract factual claims from each text segment.
Here is one example:
{Question and Segments example}
Below are your outputs:
{The Segmented output}
Below are my inputs:
{question and a list of text segments.}
@@
Table 9: Sample CoT setup prompt.
@@
I will provide you with a conversation context and the personas of the participants, that can be annotated with speaker information. As a participant in this conversation, your task is to generate a personalized response, considering the conversation context and the annotated personas. Participant Personas: [User 1 persona]: [’i attend book club every week.’ ’i am more of a cat person than a dog person.’] Conversation Context: User1: my parents dies in an accident User2: that is sad . my best friend is my mom. i could not image loosing her. User1: it is harder some days than others. but it is good you get along with your mom so well. Task Instruction: * Provide an unannotated response. * If only one persona is available, personalize the response accordingly. * If the conversation context is a single query, respond appropriately to the query. * Apply Chain of Thought reasoning to reflect on the alignment of your response with the personas. Output Format: Output a JSON of the following format: { ”reasoning”: ”briefly describe your personalization process (in 110 words or less).” ”response”: ”provide the personalized natural language response here.” }
.2 Metrics
To evaluate fluency, we use BERTScore [43] and the Naturalness evaluation metric within the UniEval framework [45]. BERTScore, implemented using the bert-score package [^14] BERTScore package: https://github.com/Tiiiger/bert_score, evaluates Precision, Recall, and F1 based on contextual embeddings, assessing how well the generated text preserves meaning and coherence relative to the reference. Nat-UniEval assesses if the response maintains human-like phrasing and tone appropriate to the dialogue history. All evaluation metrics of the UniEval framework used in our study were implemented using the official code provided in the UniEval GitHub repository [^15] https://github.com/maszhongming/UniEval. These metrics share a unified scorer function based on a fine-tuned T5 model, which calculates scores by answering Boolean-style questions specific to each evaluation dimension.
@@
For diversity assessment, we calculated Dist-1 and Dist-2 metrics [24], which respectively measure the ratio of unique unigrams and bigrams in the generated outputs. These metrics quantify lexical diversity by penalizing repetitive or generic responses, with higher scores indicating a wider vocabulary usage and more varied phrasing across generated utterances.
Table 10: Qualitative comparison of selected model responses on the FoCus dataset.
Record Prompt (User1) Model Response (excerpt) Personalization
16 “Does the homestead have any garden or lawns?”
[Persona: Loves Queensland, heritage, pastoral history]
Qwen2 7B: “Yes… beautiful landscape…”
Gemini: “Yes… impact on Aboriginal people…”
LLaMA: “… in Queensland… pastoral station…”
Qwen: Low
Gemini: Partial
LLaMA: High
450 “What happened in 1749?”
[Persona: Loves farms, architecture, dislikes rivers]
Qwen2 7B: “I’m curious too. Do you have insights?”
Gemini: “In 1749… old post offices and farms…”
LLaMA: “… settlers arrived… shows early culture…”
Qwen: Low
Gemini: Partial
LLaMA: High
858 “When was this ship decommissioned?”
[Persona: Father in Navy, lives in California, likes Panama]
Qwen2 7B: “I wonder when it was decommissioned?”
Gemini: “… April 11, 1992. Your father’s Navy background…”
LLaMA: “… decommissioned in 1992… in California, near you…”
Qwen: Low
Gemini: Partial
LLaMA: High
Personalization evaluation is conducted from two complementary perspectives: persona consistency and persona coverage, using metrics tailored to each aspect. Persona consistency reflects whether the generated response $R$ logically aligns with the given persona information $P$, while persona coverage measures how much of the persona is actually reflected or incorporated into the response. To assess consistency, we employ the C-score [32], which determines whether responses remain logically consistent with persona traits. This is achieved using an NLI model based on BERT, fine-tuned on the DNLI dataset, to predict entailment relationships between responses and persona sentences. For coverage, we use the P-dist [13], which quantifies semantic alignment between a response and persona traits. Each sentence is embedded using a pretrained encoder, and the score is computed as the minimum cosine distance between the response and the set of persona embeddings. Lower values indicate stronger semantic coverage.
To complement these personalization metrics, we also evaluate dialogue coherence using the UE-Score [23], which calculates NLI-based alignment between each utterance and its immediate response. This model, built on BERT and fine-tuned using the SNLI dataset [7], captures logical flow at the utterance level. Additionally, we incorporate the Coh-UniEval, a model-based metric that evaluates whether the response is contextually appropriate and smoothly integrated within the dialogue history. Together, these metrics provide a multifaceted evaluation of personalized dialogue systems, addressing not only surface fluency but also semantic alignment with user traits and overall conversational coherence.
As two other metrics of the UniEval framework used in the analysis diagrams, Engagingness and Groundedness are selected. Groundedness evaluates how well the response leverages and aligns with the given external factual context, ensuring relevance and accuracy. Engagingness assesses whether the response is interesting and avoids being dull or uninformative, using the dialogue history and generated responses to determine the overall engagement level.
@@
Each of these metrics offers distinct insights into the model’s performance, collectively ensuring a comprehensive and nuanced evaluation of the generated responses. When the model’s output couldn’t be evaluated because of parsing errors, we assigned default worst-case values. While discarding unparsable outputs, it is essential to ensure the automation of our automated benchmarking pipeline.
.3 Qualitative Analysis of Personalized Generation
To better illustrate the variation in personalization behavior, we provide additional examples comparing model responses across different records from the FoCus dataset. Each dialogue context includes clearly stated user persona traits, followed by a prompt. We highlight how different LLMs incorporate or overlook these traits in their final response. Table 10 summarizes three representative cases.
Across all three examples, we observe a consistent pattern: Qwen2 7B provides vague or non-committal answers, often repeating context without engaging with the user’s persona traits. This aligns with its lower personalization metric. For instance, in Record 858, Qwen2 merely re-asks the question, failing to acknowledge the user’s background. Gemini 1.5 Pro demonstrates moderate persona alignment. It correctly addresses factual content (e.g., historical context in Record 450), and occasionally references one relevant trait (e.g., “your father’s Navy background”), but does not fully weave multiple persona cues into its responses. LLaMA 3.1 8B consistently delivers highly personalized answers by grounding its response in multiple persona elements. It not only answers the prompt accurately but also enhances coherence by incorporating traits like geographical location, personal interest, and historical relevance. This demonstrates its stronger alignment with the personalization objectives in FoCus. For full dialogues, context, and persona metadata, refer to the analysis available publicly on the GitHub repository of this study  [^16] https://github.com/good-researcher/PersoBench/blob/main/response_exploration.txt.